# Parity-sensitive inhomogeneous dephasing of macroscopic spin ensembles

### v1

@@ Abstract
Spin ensembles play a pivotal role in various quantum applications such as metrology and simulating many-body physics. Recent research has proposed utilizing spin cat states to encode logical quantum information, with potentially logical lifetimes on the order of seconds via enhanced collective interactions that scale with system size. We investigate the dynamics of spin cat states under inhomogeneous broadening, revealing a phenomenon termed ‘parity-sensitive inhomogeneous dephasing’: odd cat states are significantly more susceptible to inhomogeneous dephasing compared to even cat states due to parity symmetry. Additionally, from a mean-field analysis of the driven-dissipative dynamics, we identify a synchronization phase transition wherein the ensemble becomes completely dephased beyond a critical inhomogeneous linewidth. Our findings shed light on the stability of collective spin states, important for advancing quantum technologies.
@@ I Introduction
Coherent manipulation of spin ensembles is crucial for scalable implementations of many quantum technologies, such as quantum metrology [1, 2, 3], computation [4, 5, 6] and simulation [7, 8]. In particular, spin ensembles can be engineered to behave collectively as a single high-dimensional quantum system. This allows for enhanced precision in quantum sensors [9], and is also very useful in quantum repeater protocols [10, 11]. Collective spin dynamics has also led to the discovery of interesting physical phenomena such as Dicke superradiance [12, 13], which remains an active field of study in both theoretical [14, 15, 16, 17, 18, 19, 20] and experimental [21, 22, 23] frontiers to harness it for various physical applications.
@@
In this work, we study the robustness of spin cat states in the presence of inhomogeneous broadening. This can arise for example from Doppler shifts in atomic gas clouds or spatial inhomogeneity in the electric or magnetic fields in solid state systems. Such imperfections break the permutation symmetry of the spin system, which inhibits the collective behavior. We consider the quantum driven-dissipative dynamics proposed in Ref. [27] which stabilizes the spin cat states at long times, described in Sec. II. By analytically solving for the free evolution of the spin ensemble under the sole effect of the inhomogeneous broadening in Sec. III, we uncover an effect which we term parity-sensitive inhomogeneous dephasing. We find that the robustness of the spin cat states to inhomogeneous broadening depends critically on the parity symmetry of the state, such that the even cat state is significantly more robust to inhomogeneous dephasing. This effect persists even when we consider the full quantum dynamics, suggesting that the odd cat state is indeed fragile against inhomogeneous dephasing, limiting its usefulness in encoding logical information.
In Sec. IV, we study the semiclassical mean-field dynamics derived from the full quantum model. The resulting dynamics can then be physically interpreted as a competition between spin synchronization and dephasing. We show that the system undergoes a synchronization phase transition, where the synchronization in the spin ensemble is completely broken in the long-time limit beyond a critical inhomogeneous broadening linewidth. This sets a limitation to the stability of the collective spin states against inhomogeneous dephasing. Our results provide a physical understanding of the robustness of spin cat states in realistic environments, which would be important in developing quantum technologies based on spin ensembles. We provide an outlook in Sec. V.
@@ II Driven-dissipative spin model
$\dot{\rho}=-i[H,\rho]+\frac{\Gamma_{2}}{N^{2}}\mathcal{D}\left[\sum_{i,j=1}^{N}\sigma_{i}^{-}\sigma_{j}^{-}\right]\rho$ (1)
@@
$H=\frac{1}{2}\sum_{i}\delta_{i}\sigma_{i}^{z}+\frac{\eta}{N}\left(\sum_{i,j}e^{i\varphi}\sigma_{i}^{-}\sigma_{j}^{-}+e^{-i\varphi}\sigma_{i}^{+}\sigma_{j}^{+}\right).$ (2)
@@
In the absence of inhomogeneous broadening (i.e., $\delta_{i}=0$) and the regime $\eta\ll\Gamma_{2}$, the system is weakly excited and behaves as the bosonic model studied in [28]. This can be seen by performing the Holstein-Primakoff transformation $S_{-}=(N-a^{\dagger}a)^{1/2}a$, where $S_{\pm}=\sum_{i}\sigma_{i}^{\pm}$ are the collective spin lowering and raising operators, while $a,a^{\dagger}$ are the bosonic annihilation and creation operators respectively. The spins behave collectively when $\delta_{i}=0$ due to permutation symmetry. In the weak excitation regime $\braket{a^{\dagger}a}\ll N$, we have approximately $S_{-}\approx\sqrt{N}a$ and $S_{+}\approx\sqrt{N}a^{\dagger}$. The model can be approximately described (in the rotating frame) by
@@
In Ref. [27], it was proposed to use collective spin systems in the weak excitation limit for a similar logical encoding, where the steady states are now spin coherent states [26] which can be superposed to obtain spin cat states analogous to the bosonic case. The main motivation behind this is to leverage the collective enhancement of the coherent spin interactions, such that the amplitude of the spin cat states scale as $\sim\sqrt{N}$ which translates to stronger protection against noise. It was argued that the spin cat states are robust to inhomogeneous broadening even though the permutation symmetry is broken. We now show that this depends heavily on the parity symmetry of the cat state.
@@ III Parity-sensitive inhomogeneous dephasing
Let us first consider the dynamics of the spin ensemble under the free evolution governed by the dephasing (inhomogeneous broadening) Hamiltonian $H_{0}=\sum_{n}\delta_{n}\sigma_{n}^{z}$. For concreteness, the detuning frequencies $\delta_{n}\sim\mathcal{N}(0,\delta^{2})$ are independently and identically distributed random variables drawn from a Gaussian distribution with zero mean and variance $\delta^{2}$. The overlap between the evolved state and the initial spin coherent state is given by
$\begin{split}c(t)&\equiv\braket{\theta,\phi}{e^{-iH_{0}t}}{\theta,\phi}\\ &=\prod_{n=1}^{N}e^{i\delta_{n}t/2}\left(\cos\frac{\theta}{2}+e^{-i\delta_{n}t}\sin\frac{\theta}{2}\right).\end{split}$ (6)
The fidelity $F(t)=|c(t)|^{2}$ indicates the survival probability of the initial state at time $t$. Using the Gaussian average $\mathbb{E}[\cos\delta_{n}t]=e^{-\delta^{2}t^{2}/2}$, we get the average fidelity
$\overline{F(t)}=\left[1-\frac{1}{2}(1-e^{-\delta^{2}t^{2}/2})\sin^{2}\theta\right]^{N},$ (7)
where the average is taken over random realizations of the frequencies $\delta_{n}$. In the weak excitation regime, $\theta$ is related to the amplitude of the bosonic coherent state via $|\alpha|\approx\sqrt{N}\tan(\theta/2)$ [27]. Hence, we make the small amplitude expansion $N\theta^{2}\ll 1$ to get
$\overline{F(t)}=1-\frac{N\theta^{2}}{2}(1-e^{-\delta^{2}t^{2}/2})+O(N^{2}\theta^{4})$ (8)
which implies $\overline{F(\infty)}\approx 1-(N\theta^{2})/2$ is of order unity since $N\theta^{2}\ll 1$. Using $\mathbb{E}[\cos^{2}\delta_{n}t]=(1+e^{-2\delta^{2}t^{2}})/2$, we can also evaluate the variance of the fidelity,
$\text{Var}[F(t)]=\frac{N\theta^{4}}{8}(1-e^{-\delta^{2}t^{2}})+O(\theta^{6}).$ (9)
Thus, the fluctuations of $F(t)$ are $\sim\sqrt{N}\theta^{2}$, which is much smaller than $\overline{F(t)}$. This implies that for sufficiently large systems, the random variable $F(t)$ concentrates around the mean $\overline{F(t)}$.
The spin cat states are defined as
@@
In the small amplitude regime $N\theta^{2}\ll 1$ we obtain qualitatively different results for the even cat state
$\overline{F_{+}(t)}=1-\frac{N^{2}\theta^{4}}{16}(1-e^{-\delta^{2}t^{2}})+O(N^{3}\theta^{6}),$ (14)
and the odd cat state
$\overline{F_{-}(t)}=e^{-\delta^{2}t^{2}}+O(N^{2}\theta^{4}).$ (15)
The above expressions are simplified using $N\gg 1$. Note that the leading order term in Eq. (15) should be interpreted as the limiting case of $N\theta^{2}\to 0^{+}$, since $\theta=0$ is not well defined for the odd cat state. Similar to Eq. (9), one can also show that the relative fluctuations for the fidelity vanish as $N\to\infty$.
This reveals an important fact that the robustness of the cat state against inhomogeneous broadening is very sensitive to its parity symmetry. For the even cat state, the fidelity saturates at $F_{+}\approx 1-(N\theta^{2}/4)^{2}$ which is of order unity, while for the odd cat state the fidelity vanishes rapidly on the timescale of $\delta^{-1}$. Moreover, the even cat state is more robust than the spin coherent state, since the infidelity $1-\overline{F_{+}(t)}$ is quadratic in $N\theta^{2}$ (as compared to linear for the spin coherent state). Physically, this means that the dephasing effects are suppressed by even symmetry and amplified by odd symmetry.
III.1 Adding the stabilization dynamics

### v2

# Stability of macroscopic spin ensembles against inhomogeneous dephasing

@@ Abstract
Spin ensembles play a pivotal role in various quantum applications such as metrology and simulating many-body physics. Recent research has proposed utilizing spin cat states to encode logical quantum information, with logical lifetimes potentially on the order of seconds, achieved via enhanced collective interactions that scale with system size. We investigate the dynamics of spin cat states under inhomogeneous broadening, revealing a phenomenon termed ‘parity-sensitive inhomogeneous dephasing’: for small amplitudes, odd cat states are significantly more susceptible to inhomogeneous dephasing compared to even cat states due to the difference in parity symmetry. This discrepancy between even and odd cat states vanishes at large amplitudes, and behave similarly to a spin coherent state with the same amplitude. To analyze the stability of the spin coherent state, we perform a mean-field analysis of the driven-dissipative dynamics, from which we identify a synchronization phase transition wherein the ensemble becomes completely dephased beyond a critical inhomogeneous linewidth. The mean-field analysis suggests that the dissipative stabilization can suppress the decoherence effects from inhomogeneous broadening. We argue that the stability of the mean-field model provides a reasonable estimate for that of spin cat states with a large amplitude in the full quantum model. Our findings shed light on the stability of collective spin states, important for advancing quantum technologies.
@@ I Introduction
Coherent manipulation of spin ensembles is crucial for scalable implementations of many quantum technologies, such as quantum metrology [1, 2, 3], computation [4, 5, 6] and simulation [7, 8]. In particular, spin ensembles can be engineered to behave collectively as a single high-dimensional quantum system. This allows for enhanced precision in quantum sensors [9], and is also very useful in quantum repeater protocols [10, 11]. Collective spin dynamics has also led to the discovery of interesting physical phenomena such as Dicke superradiance [12, 13], which remains an active field of study in both theoretical [14, 15, 16, 17, 18, 19, 20, 21] and experimental [22, 23, 24] frontiers to harness it for various physical applications.
@@
In this work, we study the robustness of spin cat states in the presence of inhomogeneous broadening. This can arise for example from Doppler shifts in atomic gas clouds or spatial inhomogeneity in the electric or magnetic fields in solid state systems. Such imperfections break the permutation symmetry of the spin system, which inhibits the collective behavior. We consider the quantum driven-dissipative dynamics proposed in Ref. [28] which stabilizes the spin cat states at long times, described in Sec. II. The central theme of the paper is to address the following question:
In large ensembles of spins, can spin cat states with dissipative stabilization be robust against inhomogeneous broadening?
By analytically solving for the free evolution of the spin ensemble under the sole effect of the inhomogeneous broadening in Sec. III, we uncover an effect which we term parity-sensitive inhomogeneous dephasing. We find that the robustness of the spin cat states to inhomogeneous broadening at small amplitudes depends critically on the parity symmetry of the state, such that the even cat state is significantly more robust to inhomogeneous dephasing than the odd cat state. At large amplitudes, the discrepancy between even and odd cat states vanishes, and the robustness of cat states are similar to that of a spin coherent state with the same amplitude.
Motivated by this observation, we focus on the robustness of the spin coherent state, in the presence of dissipative stabilization. To this end, we study in Sec. IV the semiclassical mean-field dynamics derived from the full quantum model. The resulting dynamics can then be physically interpreted as a competition between spin synchronization and dephasing. We show that the system undergoes a synchronization phase transition, where the synchronization in the spin ensemble is completely broken in the long-time limit beyond a critical inhomogeneous broadening linewidth. This sets a limitation to the stability of the collective spin states against inhomogeneous dephasing. While the mean-field model is, by construction, inadequate in describing quantum effects, we argue that the analysis provides useful insights into the stability of spin cat states with large amplitudes in the full quantum model. Importantly, our findings suggest that in a realistic implementation of the quantum model, the spin cat state of either parity at large amplitudes is robust under the condition that the broadening linewidth does not scale with the system size. Our results provide a physical understanding of the robustness of spin cat states in realistic environments, which would be important in developing quantum technologies based on spin ensembles. We provide an outlook in Sec. V.
@@ II Driven-dissipative spin model
$\dot{\rho}=-i[H,\rho]+\frac{\Gamma_{2}}{N^{2}}\mathcal{D}\left[\sum_{m,n=1}^{N}\sigma_{m}^{-}\sigma_{n}^{-}\right]\rho$ (1)
@@
$H=\frac{1}{2}\sum_{n=1}^{N}\delta_{n}\sigma_{n}^{z}+\frac{\eta}{N}\left(\sum_{m,n=1}^{N}e^{i\varphi}\sigma_{m}^{-}\sigma_{n}^{-}+e^{-i\varphi}\sigma_{m}^{+}\sigma_{n}^{+}\right).$ (2)
@@
In the absence of inhomogeneous broadening (i.e., $\delta_{n}=0$) and the regime $\eta\ll\Gamma_{2}$, the system is weakly excited and behaves as the bosonic model studied in [29]. This can be seen by performing the Holstein-Primakoff transformation $S_{-}=(N-a^{\dagger}a)^{1/2}a$, where $S_{\pm}=\sum_{n}\sigma_{n}^{\pm}$ are the collective spin lowering and raising operators, while $a,a^{\dagger}$ are the bosonic annihilation and creation operators respectively. The spins behave collectively when $\delta_{n}=0$ due to permutation symmetry. In the weak excitation regime $\braket{a^{\dagger}a}\ll N$, we have approximately $S_{-}\approx\sqrt{N}a$ and $S_{+}\approx\sqrt{N}a^{\dagger}$. The model can be approximately described (in the rotating frame) by
@@
In Ref. [28], it was proposed to use collective spin systems in the weak excitation limit for a similar logical encoding, where the steady states are now spin coherent states [27] which can be superposed to obtain spin cat states analogous to the bosonic case. The main motivation behind this is to leverage the collective enhancement of the coherent spin interactions, such that the amplitude of the spin cat states scale as $\sim\sqrt{N}$ which translates to stronger protection against noise. It was argued that the spin cat states are robust to inhomogeneous broadening even though the permutation symmetry is broken. We now show that, for small amplitudes, this depends heavily on the parity symmetry of the cat state.
@@ III Parity-sensitive inhomogeneous dephasing
By considering even and odd superpositions of |$\ket{\theta,\phi}$ and |$\ket{\theta,\phi+\pi}$, one obtains the even and odd spin cat states
@@
|$\ket{\text{Cat}_{\pm}}$ are eigenstates of the parity operator
$\hat{\Pi}=\exp\left(i\pi\sum_{n}\sigma_{n}^{+}\sigma_{n}^{-}\right)$ (8)
with eigenvalues ±$\pm 1$ respectively. Thus, |$\ket{\text{Cat}_{\pm}}$ has even/odd number of excitations. Similar to the bosonic cat states, spin cat states are potentially useful for encoding a logical qubit and in quantum metrology. For a sufficiently small inhomogeneous broadening, the dissipative model (1) approximately encodes arbitrary superpositions of the spin coherent states in the steady state manifold. Thus, the spin cat states |$\ket{\text{Cat}_{\pm}}$ can be prepared by initializing the spins in a state of definite parity (i.e., an eigenstate of $\hat{\Pi}$). For example, by initializing the spins in the ground state (corresponding to the vacuum state in the bosonic picture), the steady state is approximately |$\ket{\text{Cat}_{+}}$. The validity of this approximation for the steady state depends on the stability of the cat states under inhomogeneous broadening, which is the core focus of this paper.
In this section, we study the robustness of the spin cat states against inhomogeneous broadening. The main result here is that for small amplitudes $|\alpha|\ll 1$, the dephasing effect is sensitive to the parity of the cat state, with |$\ket{\text{Cat}_{+}}$ being significantly more robust than |$\ket{\text{Cat}_{-}}$. This discrepancy vanishes at large amplitudes $|\alpha|\gg 1$, and both |$\ket{\text{Cat}_{\pm}}$ experience similar dephasing effects.
III.1 Free evolution with inhomogeneous dephasing
Let us first consider the dynamics of the spin ensemble under the free evolution governed by the dephasing (inhomogeneous broadening) Hamiltonian $H_{0}=\sum_{n}\delta_{n}\sigma_{n}^{z}/2$. For concreteness, the detuning frequencies $\delta_{n}\sim\mathcal{N}(0,\delta^{2})$ are independently and identically distributed random variables drawn from a Gaussian distribution with zero mean and variance $\delta^{2}$. The overlap between the evolved state and the initial spin coherent state is given by
$\begin{split}c(t)&\equiv\braket{\theta,\phi}{e^{-iH_{0}t}}{\theta,\phi}\\ &=\prod_{n=1}^{N}e^{i\delta_{n}t/2}\left(\cos^{2}\frac{\theta}{2}+e^{-i\delta_{n}t}\sin^{2}\frac{\theta}{2}\right).\end{split}$ (9)
The fidelity $F(t)=|c(t)|^{2}$ indicates the survival probability of the initial state at time $t$, which evaluates to
$\begin{split}F(t)&=\prod_{n=1}^{N}\left|\cos^{2}\frac{\theta}{2}+e^{-i\delta_{n}t}\sin^{2}\frac{\theta}{2}\right|^{2}\\ &=\prod_{n=1}^{N}\left[\cos^{4}\frac{\theta}{2}+\sin^{4}\frac{\theta}{2}+2\cos(\delta_{n}t)\sin^{2}\frac{\theta}{2}\cos^{2}\frac{\theta}{2}\right].\end{split}$ (10)
Using the Gaussian average $\mathbb{E}[\cos\delta_{n}t]=e^{-\delta^{2}t^{2}/2}$, we get the disorder-averaged fidelity
$\begin{split}\overline{F(t)}&=\left[\cos^{4}\frac{\theta}{2}+\sin^{4}\frac{\theta}{2}+2e^{-\delta^{2}t^{2}/2}\sin^{2}\frac{\theta}{2}\cos^{2}\frac{\theta}{2}\right]^{N}\\ &=\left[1-2\left(1-e^{-\delta^{2}t^{2}/2}\right)\sin^{2}\frac{\theta}{2}\cos^{2}\frac{\theta}{2}\right]^{N}\\ &=\left[1-\frac{1}{2}\left(1-e^{-\delta^{2}t^{2}/2}\right)\sin^{2}\theta\right]^{N},\end{split}$ (11)
where the average is taken over random realizations of the frequencies $\delta_{n}$. In the weak excitation regime, $\theta$ is related to the amplitude of the bosonic coherent state via $|\alpha|\approx\sqrt{N}\tan(\theta/2)$ [28]. We will work in the regime where the bosonic amplitude $|\alpha|=\sqrt{2\eta/\Gamma_{2}}\ll\sqrt{N}$ such that $\theta\approx 2|\alpha|/\sqrt{N}\ll 1$. Note that $|\alpha|$ can still be much greater than $1$ for a sufficiently large $N$. Hence,
$\overline{F(t)}\approx\exp\left(-2|\alpha|^{2}(1-e^{-\delta^{2}t^{2}/2})\right).$ (12)
For small amplitudes $|\alpha|\ll 1$,
$\overline{F(t)}=1-2|\alpha|^{2}(1-e^{-\delta^{2}t^{2}/2})+O(|\alpha|^{4})$ (13)
which implies $\overline{F(\infty)}\approx 1-2|\alpha|^{2}$ is of order unity. Using $\mathbb{E}[\cos^{2}\delta_{n}t]=(1+e^{-2\delta^{2}t^{2}})/2$, we can also evaluate the variance of the fidelity,
$\text{Var}[F(t)]=\frac{2|\alpha|^{4}}{N}(1-e^{-\delta^{2}t^{2}})+O(|\alpha|^{6}).$ (14)
Thus, the fluctuations of $F(t)$ are $\sim|\alpha|^{2}/\sqrt{N}$, much smaller than $\overline{F(t)}$. This implies that for sufficiently large systems, the random variable $F(t)$ concentrates around the mean $\overline{F(t)}$.
@@ III.1 Free evolution with inhomogeneous dephasing
Using the approximation
$\cos^{N}{\theta}\approx\left(1-\frac{2|\alpha|^{2}}{N}\right)^{N}\approx e^{-2|\alpha|^{2}},$ (17)
the fidelity is expressed as
$\begin{split}\overline{F_{\pm}(t)}&\approx\frac{1}{(1\pm e^{-2|\alpha|^{2}})^{2}}\bigg{[}\exp\left(-2|\alpha|^{2}(1-e^{-\delta^{2}t^{2}/2})\right)\\ &+\exp\left(-2|\alpha|^{2}(1+e^{-\delta^{2}t^{2}/2})\right)\pm 2e^{-2|\alpha|^{2}}\bigg{]}.\end{split}$ (18)
In the long-time limit $t\to\infty$, we have
$\begin{split}\overline{F_{\pm}(\infty)}&\approx\frac{2}{(1\pm e^{-2|\alpha|^{2}})^{2}}\left(e^{-2|\alpha|^{2}}\pm e^{-2|\alpha|^{2}}\right)\\ &=\begin{cases}\operatorname{sech}^{2}(|\alpha|^{2})\quad&(\text{even cat}),\\ 0\quad&(\text{odd cat}).\end{cases}\end{split}$ (19)
This shows that |$\ket{\text{Cat}_{-}}$ always decays to zero. On the other hand, the long-time fidelity for |$\ket{\text{Cat}_{+}}$ depends on the amplitude. For small amplitude $|\alpha|\ll 1$, $\overline{F_{+}(\infty)}=1-|\alpha|^{4}+O(|\alpha|^{8})$ is robust, while for large amplitude $|\alpha|\gg 1$, $\overline{F_{+}(\infty)}\approx 4e^{-2|\alpha|^{2}}$ vanishes exponentially with $|\alpha|$.
Next, we look at the short-time behavior of $\overline{F_{\pm}(t)}$. Expanding Eq. (18) to order $t^{2}$ gives
$\overline{F_{\pm}(t)}\approx\begin{cases}1-|\alpha|^{2}\tanh(|\alpha|^{2})\delta^{2}t^{2}\quad&\text{(even cat)},\\ 1-|\alpha|^{2}\coth(|\alpha|^{2})\delta^{2}t^{2}\quad&\text{(odd cat)}.\end{cases}$ (20)
For small amplitude $|\alpha|\ll 1$,
$\overline{F_{\pm}(t)}\approx\begin{cases}1-|\alpha|^{4}\delta^{2}t^{2}\quad&\text{(even cat)},\\ 1-\delta^{2}t^{2}\quad&\text{(odd cat)},\end{cases}$ (21)
implying that |$\ket{\text{Cat}_{-}}$ decays faster than |$\ket{\text{Cat}_{+}}$. For large amplitudes $|\alpha|\gg 1$, $\overline{F_{\pm}(t)}\approx 1-|\alpha|^{2}\delta^{2}t^{2}$ and both |$\ket{\text{Cat}_{\pm}}$ decay identically, and similarly as the spin coherent state. As with Eq. (14), one can also show that the relative fluctuations for the fidelity vanish as $N\to\infty$.
This reveals an important fact that the robustness of the cat state against inhomogeneous broadening is very sensitive to its parity symmetry at small amplitudes. For the even cat state, the fidelity saturates at $F_{+}\approx 1-|\alpha|^{2}$ which is of order unity, while for the odd cat state the fidelity vanishes rapidly on the timescale of $\delta^{-1}$. Moreover, the even cat state is more robust than the spin coherent state, since the infidelity $1-\overline{F_{+}(t)}\sim|\alpha|^{4}$ (as compared to $|\alpha|^{2}$ for the spin coherent state). Physically, this means that the dephasing effects are suppressed by even symmetry and amplified by odd symmetry. The parity-sensitive dephasing of the cat states, at small amplitudes $|\alpha|\ll 1$, can be understood as an interference effect, where constructive and destructive interference occur for the even and odd cat states respectively. The strength of the interference is given by $|\braket{\theta,\phi}{\theta,\phi+\pi}|^{2}=\cos^{4N}\theta$, which appears in the cross terms of Eq. (15) upon expanding |$\ket{\text{Cat}_{\pm}}$ in terms of the spin coherent states. At sufficiently large amplitudes, the interference strength is close to zero (as $N\to\infty)$, which is equivalent to the fact that |$\ket{\theta,\phi}$ and |$\ket{\theta,\phi+\pi}$ become nearly orthogonal. Consequently, the discrepancy in the fidelity decay between spin cat states of different parities vanishes at large amplitudes.
III.2 Adding the stabilization dynamics